# Trapped-ion Laser Cooling in Structured Light Fields

### v2

@@ III Conclusion
By selectively driving desired sideband transitions, carrier-free cooling at nodal points of SW or HG modes allows advantages over cooling with running waves in cooling speed and final phonon limit. Modest advantages obtain for Doppler cooling of highly excited ions from well beyond the LDR, with significant enhancements predicted for GS cooling via EIT including in motional mode frequency bandwidth addressed. The proposed schemes utilize either simple SW or HG modes that can be stably and robustly delivered to ions in integrated photonic architectures. Our results indicate the potential for scalable photonics to assist in bypassing key physical limitations of trapped-ion systems, and inform experimental realizations in progress.
@@ IV Acknowledgments
Appendix A Lindblad master equation beyond the LDR
The dynamics of a trapped ion in a harmonic potential, interacting with a laser field, can be described using the Lindblad master equation. Inside LDR, Eq. (1) is approximated to first order $\langle e|V_{dip}|g\rangle\approx\frac{\Omega_{0}}{2}(1+i\eta(\hat{a}+\hat{a}^{\dagger}))$, resulting in a carrier and two first-order sideband transitions. The strongest present coupling, carrier transition, governs the internal states of the ion. This allows us to approximate the system by decoupling the external motional states from the internal dynamics and describe the cooling behavior with the rate equation Cirac et al. (1992); Stenholm (1986). These approximations fail when going outside the LDR or when sideband transitions become the leading order coupling as in the carrier-free cooling configurations considered here. Thus, to investigate the dynamics outside the LDR and quantify the enhancement limit of carrier-free cooling, we directly simulate the Lindblad master equation with higher-order terms.
Figure 6: Comparison of simulation (points) and analytical results (lines) for cooling rate over $\omega_{m}$ for the low-intensity limit of cooling beam. The pump beam’s intensity and cooling beam’s detuning are optimized for $\omega_{m0}=2\pi\times 3$ MHz and $\Delta_{p}=5\Gamma$. Blue and red represent results for cooling beam delivered as a SW and RW, respectively. (a) uses a SW pump beam with the ion at an antinode, with $\eta_{c}=\eta_{0}$ and $\Omega_{c}=2\pi\times 2$ MHz for both SW and RW cooling beam. (b) uses RW pump beam. Limited by the selection rule, we choose $\eta_{c}=-\eta_{p}=\eta_{0}/\sqrt{2}$ for RW EIT to maximized $(\eta_{c}-\eta_{p})^{2}$ as in the usual experimental design Lechner et al. (2016). For carrier-free EIT, we choose $\eta_{p}=\eta_{0}/\sqrt{2}$ for the RW pump and $\eta_{c}=\eta_{0}$ for the SW cooling beam for simplicity as only the coupling term $\eta_{c}\Omega_{c}$ matters at SW node in the LDR. The same LD parameters are used in the simulations shown in previous sections. Lastly, we choose $\Omega_{c}=2\pi\times 2\sqrt{2}$ MHz for SW cooling beam and $\Omega_{c}=2\pi\times 2$ MHz for RW cooling beam so that their cooling rates at $\omega_{m0}$ agree in QRT calculation. The presence of pump-beam sideband couplings present in (b) distorts the good agreement between the analytical and master-equation treatment in (a); both treatments show the enhancement in motional frequency bandwidth associated with SW cooling beam delivery.
@@ Appendix A Lindblad master equation beyond the LDR
with $\hat{A}=e^{ixk_{0}\hat{R}}\hat{\sigma}^{-},\hat{\sigma}^{-}=|g\rangle\langle e|$, $x=\cos(\theta)$ for emission angle $\theta$ with respect to motional axis Cirac et al. (1992). Using the Baker–Campbell–Hausdorff formula, $e^{ixk_{0}\hat{R}}=e^{ix\eta_{0}(\hat{a}^{{\dagger}}+\hat{a})}=e^{-\eta_{0}^{2}x^{2}/2}e^{ix\eta_{0}\hat{a}^{{\dagger}}}e^{ix\eta_{0}\hat{a}}$. Denote $\hat{b}_{k}=\frac{(i\eta_{0}\hat{a}^{{\dagger}})^{k}}{k!},\hat{d}_{k}=\frac{(i\eta_{0}\hat{a})^{k}}{k!}$ ($\hat{b}_{j}=\hat{d}_{j}=0\text{ if }j<0\text{ or }j>n$), and Taylor expanding $\hat{A}$ to the $n$th term, we can then calculate the jump operator for the Liouvillian with ±mplus-or-$\pm m$ ($0\leq m\leq n$) change in motional states as
@@ Appendix B Analytical EIT treatment
Our analytical description of carrier-free EIT cooling follows the approach of Zhang et al. (2012) for an ideal three-level system. A three-level system withpump beam of carrier strength $\Omega_{p0}$ and sideband strength $\eta_{p}\Omega_{p1}$ (coupling states $|r\rangle$ and $|e\rangle$) and cooling beam of carrier strength $\Omega_{c0}$ and sideband strength $\eta_{c}\Omega_{c1}$ (coupling states $|g\rangle$ and $|e\rangle$) can be represented by a rotating-frame Hamiltonian:
@@
where $\mathcal{F}$ denotes the Fourier transform with respect to $\tau$. Using the quantum regression theorem (QRT) Gardiner and Zoller (2004), we can solve for an analytical expression of (5) from the steady-state solution to $\hat{H}_{\text{int}}$.
@@
$\text{Re}[\mathcal{S}(\nu)]=\frac{2\eta_{c}^{2}\Omega_{c1}^{2}\Gamma(\Delta+\nu)^{2}}{4\Gamma^{2}(\Delta+\nu)^{2}+\big{(}\Omega_{p}^{2}-4(\Delta+\nu)(\Delta_{c}+\nu)\big{)}^{2}},$
where $\Gamma=\Gamma_{c}+\Gamma_{p}$ and $\Delta=\Delta_{c}-\Delta_{p}$. When $\Delta=\nu$, this gives the cooling rate in Zhang et al. (2012) for ideal carrier-free EIT, which does not depend on $\eta_{p}$.
For the standard RW EIT cooling, taking the leading order of $\Omega_{c}$ again, (5) reduces to a simple expression when $\Delta=\Delta_{c}-\Delta_{p}$, proportional to $(\eta_{c}-\eta_{p})^{2}\Omega_{c1}^{2}$. This motivates the counterpropagating cooling and pump beam in experiments Lechner et al. (2016).
The analytical treatment above assumes that sideband couplings do not affect the steady state solution to the optical Bloch equation, valid in the LDR and in the low intensity limit for both $\Omega_{c1}$ and $\Omega_{p1}$. In reality, the effect of the pump beam’s sideband coupling on the internal dynamics is always nonnegligible. Fig. 6 shows cooling rates for for EIT using SW and RW pump beam, comparing the simulation with the analytical results from (5) under the low-intensity limit of cooling beam. For the SW pump beam, we place the ion at its anti-node, where only the carrier is present. This discrepancy between the simulation and analytical results for RW pump beam showcases the necessity of full master simulation in understanding the physical cooling performance.

### v3

@@ III Conclusion
By selectively driving desired sideband transitions, carrier-free cooling at nodal points of SW or HG modes allows advantages over cooling with running waves in cooling speed and final phonon number limit. Modest advantages obtain for Doppler cooling of highly excited ions from well beyond the LD regime, with significant enhancements predicted for ground-state cooling via EIT including in motional mode frequency bandwidth addressed. The proposed schemes utilize either simple SW or HG modes that can be stably and robustly delivered to ions in integrated photonic architectures. Furthermore, carrier nulling may also allow resolved sideband cooling at high rates that with RW fields would be limited by carrier excitation due to power broadening. Our results indicate potential for scalable photonics to assist in alleviating key physical limitations of trapped-ion systems, and inform experimental realizations in progress.
@@ IV Acknowledgments
V Data availability
The data that support the findings of this article are openly available [55].
Appendix A Lindblad master equation beyond the LD regime
The dynamics of a trapped ion in a harmonic potential, interacting with a laser field, can be described using the Lindblad master equation. Inside LD regime, Eq. (1) is approximated to first order $\langle e|V_{dip}|g\rangle\approx\frac{\Omega_{0}}{2}(1+i\eta(\hat{a}+\hat{a}^{\dagger}))$, resulting in a carrier and two first-order sideband transitions. The strongest present coupling, carrier transition, governs the internal states of the ion. This allows us to approximate the system by decoupling the external motional states from the internal dynamics and describe the cooling behavior with the rate equation [25, 23]. These approximations fail when going outside the LD regime or when sideband transitions become the leading order coupling as in the carrier-free cooling configurations considered here. Thus, to investigate the dynamics outside the LD regime and quantify the enhancement limit of carrier-free cooling, we directly simulate the Lindblad master equation with higher-order terms using a Monte-Carlo method.
@@ Appendix A Lindblad master equation beyond the LD regime
with $\hat{A}=e^{ixk_{0}\hat{R}}\hat{\sigma}^{-},\hat{\sigma}^{-}=|g\rangle\langle e|$, $x=\cos(\theta)$ for emission angle $\theta$ with respect to motional axis [25]. Using the Baker–Campbell–Hausdorff formula, $e^{ixk_{0}\hat{R}}=e^{ix\eta_{0}(\hat{a}^{{\dagger}}+\hat{a})}=e^{-\eta_{0}^{2}x^{2}/2}e^{ix\eta_{0}\hat{a}^{{\dagger}}}e^{ix\eta_{0}\hat{a}}$. Denote $\hat{b}_{k}=\frac{(i\eta_{0}\hat{a}^{{\dagger}})^{k}}{k!},\hat{d}_{k}=\frac{(i\eta_{0}\hat{a})^{k}}{k!}$ ($\hat{b}_{j}=\hat{d}_{j}=0\text{ if }j<0\text{ or }j>n$), and Taylor expanding $\hat{A}$ to the $n$th term, we can then calculate the jump operator in the Liouvillian with ±mplus-or-$\pm m$ ($0\leq m\leq n$) change in motional states as
@@ Appendix B Analytical EIT treatment
Figure 6: Comparison of simulation (points) and analytical results (lines) for cooling rate over $\omega_{m}$ for the low-intensity limit of cooling beam. The pump beam’s intensity and cooling beam’s detuning are optimized for $\omega_{m0}=2\pi\times 3$ MHz and $\Delta_{p}=5\Gamma$. Blue and red represent results for cooling beam delivered as a SW and RW, respectively. (a) uses a SW pump beam with the ion at an antinode, with $\eta_{c}=\eta_{0}$ and $\Omega_{c}=2\pi\times 2$ MHz for both SW and RW cooling beam. (b) uses RW pump beam. We choose $\eta_{c}=-\eta_{p}=\eta_{0}/\sqrt{2}$ for RW EIT and $\eta_{c}=\eta_{0}$ for the SW cooling beam as discussed in the previous section. The same LD parameters are used in the simulations shown in previous sections. Lastly, we choose $\Omega_{c}=2\pi\times 2\sqrt{2}$ MHz for SW cooling beam and $\Omega_{c}=2\pi\times 2$ MHz for RW cooling beam so that their cooling rates at $\omega_{m0}$ agree in quantum regression theorem calculation. The presence of pump-beam sideband couplings present in (b) distorts the good agreement between the analytical and master-equation treatment in (a); both treatments show the enhancement in motional frequency bandwidth associated with SW cooling beam delivery.
Our analytical description of carrier-free EIT cooling follows the approach of [26] for an ideal three-level system. A three-level system with pump beam of carrier strength $\Omega_{p0}$ and sideband strength $\eta_{p}\Omega_{p1}$ (coupling states $|r\rangle$ and $|e\rangle$) and cooling beam of carrier strength $\Omega_{c0}$ and sideband strength $\eta_{c}\Omega_{c1}$ (coupling states $|g\rangle$ and $|e\rangle$) can be represented by a rotating-frame Hamiltonian:
@@
where $\mathcal{F}$ denotes the Fourier transform with respect to $\tau$. Using the quantum regression theorem [56], we can solve for an analytical expression of (6) from the steady-state solution to $\hat{H}_{\text{int}}$.
@@
$\text{Re}[\mathcal{S}(\nu)]=\frac{2\eta_{c}^{2}\Omega_{c1}^{2}\Gamma(\Delta+\nu)^{2}}{4\Gamma^{2}(\Delta+\nu)^{2}+\big{(}\Omega_{p0}^{2}-4(\Delta+\nu)(\Delta_{c}+\nu)\big{)}^{2}},$
where $\Gamma=\Gamma_{c}+\Gamma_{p}$ and $\Delta=\Delta_{c}-\Delta_{p}$. When $\Delta=\nu$, this gives the cooling rate in [26] for ideal carrier-free EIT, which does not depend on $\eta_{p}$.
For the standard RW EIT cooling with $\Omega_{c0}=\Omega_{c1}$ and $\Omega_{p0}=\Omega_{p1}$, taking the leading order of $\Omega_{c}$ again, (6) reduces to a simple expression when $\Delta=\Delta_{c}-\Delta_{p}$, proportional to $(\eta_{c}-\eta_{p})^{2}\Omega_{c1}^{2}$.
The analytical treatment above assumes that sideband couplings do not affect the stead-state solution to the optical Bloch equation, valid in the LD regime and in the low intensity limit for both $\Omega_{c1}$ and $\Omega_{p1}$. In reality, the effect of the pump beam’s sideband coupling on the internal dynamics is always non-negligible. Fig. 6 shows cooling rates for EIT using SW and RW pump beam, comparing the simulation with the analytical results from (6) under the low-intensity limit of cooling beam. For the SW pump beam, we consider an ion positioned at the SW’s antinode where only the carrier is present. The discrepancy between the simulation and analytical results for RW pump beam indicates the need for full master equation simulation.
Appendix C Position Deviation
The carrier-free cooling schemes depend strongly on the suppression of carrier transition by placing the ion at the intensity null. Thus, the effect of ion’s position inside the SW is discussed previously for both Doppler and EIT cooling [25, 26]. To complete the discussion, we simulate the cooling dynamics for the eight-level ${}^{40}\text{Ca}^{+}$ at different $\phi$ inside the SW $\sin(\bm{k}\cdot\hat{\bm{R}}+\phi)$ in Fig. 7. The results suggest that when the ion is positioned within $|\phi|\lesssim\frac{\pi}{16}$, the final EIT cooling limit remains within one order of magnitude of that at the SW node. For Doppler cooling, the same positional tolerance results in negligible impact on final phonon number. For purely counterpropagating SW beams, this corresponds to a positional deviation of $|\Delta x|\lesssim\frac{\lambda}{32}$, i.e. a positional accuracy of within approximately ±$\pm 10$ nm. Previous observations of trapped-ion control in integrated photonic devices indicate the practicality of such positioning [28]. We note further that for SW beams intersecting at an angle (i.e. with a lower wavevector projection along the motional mode axis) the positional accuracy required to achieve $\phi\lesssim\frac{\pi}{16}$ is further relaxed.
Figure 7: Sensitivity of both SW Doppler and EIT cooling to positional offset from the SW node, for $\omega_{m}=2\pi\times 3$ MHz. Positional offset is defined in terms of phase offset $\phi$ within the SW field with spatial dependence $\sin(\bm{k}\cdot\hat{\bm{R}}+\phi)$. (a) Doppler cooling limit with $\Delta_{c}=-0.5\Gamma$ and $\Omega_{c}=1.8\Gamma$ as in Fig. 2b. (b) EIT cooling rate and limit with optimal parameters for $\omega_{m}=2\pi\times 3$ MHz, $\Delta_{p}=5\Gamma$ and $\Omega_{c}=2\pi\times 42$ MHz as in Fig. 5a.
Appendix D Motional bandwidth for carrier-free Doppler cooling
Figure 8: Doppler cooling limit in RW (left) and SW (right) fields as a function of $\Omega_{c}$ and $\omega_{m}$ with $\Delta_{c}=-\Gamma/2$, calculated from the eight-level simulation.
In Fig. 4, we observe that carrier-free EIT when optimized for a center motional frequency can cool a larger motional bandwidth as compared to RW EIT. For Doppler cooling, both RW and carrier-free cooling should be able to cool a broad range of motional frequencies. Within the LD regime and for low cooling beam amplitude $\Omega_{c}$ set to $\Delta_{c}=-\Gamma/2$, we expect $\bar{n}_{ss}=\frac{\Gamma}{2\omega_{m}}$ for RW and $\bar{n}_{ss}=\frac{\Gamma}{4\omega_{m}}$ for carrier-free Doppler cooling [25]. This behavior is apparent in the simulation results for $\bar{n}_{\mathrm{ss}}$ from the eight-level simulation, particularly at large $\Omega_{m}$ for which the LD limit applies (Fig. 8).